solve and check the quadratic equation below by taking the square root of both sides. express irrational answers in radical form.
(x-10)^2=4
check

Answer by stanbon(75887)

solve and check the quadratic equation below by taking the square root of both sides. express irrational answers in radical form.
------------
(x-10)^2=4
---
x-10 = 2 or x-10 = -2
x = 12 or x = 8
======================
Cheers,
Stan H.